A wave has a wave speed of 243 m/s and a wavelength of 3.27 cm. Calculate (a) the frequency and (b) the period of the wave.
 

Grade:12th pass

1 Answers

Vikas TU
4 years ago
Dear student 
Velocity=frequency × wavelength
v= f×λ
By putting the values
243m/s=0.0327m×frequency
Frequency= 243÷0.0327
Frequency≈7431.19
Time period =1/frequency
T = 1/7431.19 
= 1.34* 10^-4